I need step by step on solving
LOG 10/3 without a calculator
This is what I tried:
log 10/3 = x
10^x = 10/3
How do I solve this problem step by step from this point?

Answer by CharlesG2(834)

I need step by step on solving
LOG 10/3 without a calculator
This is what I tried:
log 10/3 = x
10^x = 10/3
How do I solve this problem step by step from this point?
LOG RULES
1) logb(mn) = logb(m) + logb(n)
2) logb(m/n) = logb(m) – logb(n)
3) logb(mn) = n · logb(m)
y = b^x is same thing as logb(y) = x
log10 (10/3) = x is sane thing as 10^x=(10/3)
using rule 2
log10 (10/3) = log10 (10) - log10 (3) = x
1 - log10 (3) = x